# Integrating Various Software Artifacts for Better LLM-based Bug Localization and Program Repair

### v2

@@ 6.2. Threats to Validity
Another potential threat is that the ChatGPT-4o-mini model used in this work may have been trained on open-source projects from GitHub, which could overlap with the Defects4J dataset, leading to possible data leakage. To mitigate this, we also randomly selected 100 bugs from another dataset, GrowingBugs (Jiang et al., 2022), and found the plausible fixing rate to be 39%, which may help alleviate this concern. Additionally, the debugging information requires dynamic analysis, which is unlikely to have been used during the model’s training. Also as found in (Li et al., 2024), directly using the GPT-4 model can not improve the fix rate.
The final threat to the external validity is that our experiments were conducted on Java projects, and the findings may not be generalized to projects written in other programming languages. To address this, we plan to design MethodRecorder and DebugRecorder on other programming languages and evaluate DEVLoRe on additional datasets across multiple programming languages in our future work.
@@ 7.1. Large Language Models for Fault Localization
Recently, there has been significant interest in using LLMs for fault localization. Toggle incorporated additional contextual information, such as the buggy line number or code review comments, and greatly enhances the accuracy of predicting both the starting and ending buggy tokens (Hossain et al., 2024). AGENTFL employs a multi-agent system based on ChatGPT and frames the fault localization task as a three-step process: comprehension, navigation, and confirmation. In each step, AGENTFL deploys specialized agents, each with unique expertise, and uses different tools to address specific tasks (Qin et al., 2024). CrashTracker conducts static analysis to map each crash to the corresponding exception instance and identify potential buggy candidates. It then utilizes LLMs to enhance the explainability of the localization results (Yan et al., 2024). Jiang et al. assessed the performance of recent commercial closed-source general-purpose LLMs, such as ChatGPT 3.5, ERNIE Bot 3.5, and IFlytek Spark 2.0, on line-level fault localization with the provided buggy method (Jiang et al., 2024). LLMAO fine-tunes LLMs with 350M, 6B, and 16B parameters on small, curated corpora like Defects4J, improving Top-1 fault localization by 2.3%-54.4% and Top-5 results by 14.4%-35.6%, compared to the state-of-the-art machine learning fault localization (Yang et al., 2024a). AutoFL prompts an LLM to use function calls for navigating a repository, enabling effective fault localization in large codebases while overcoming the LLM context length limit. It also generates an explanation of the bug and suggests a fault location (Kang et al., 2024). LLM4FL combines traditional spectrum-based fault localization with prompt chaining to divide large coverage data into manageable groups. By employing multiple LLM agents, it navigates the codebase more effectively to localize faults (Rafi et al., 2024). Like LLM4FL and AutoFL which combines transitional fault localization tools, DEVLoRe incorporates a static code analysis tool, MethodRecorder, to identify relevant buggy methods in a lightweight manner by invoking failing tests. However, unlike the above approaches, we provide the LLM with a combination of software artifacts, including issue content, error stack traces, and debugging information, which are commonly used by human developers for fault localization. This enables DEVLoRe to outperform state-of-the-art fault localization methods with greater efficiency and lower cost.
@@ 7.2. Large Language Models for Program Repair
Recent studies have extensively explored the use of LLMs for program repair. ChatRepair initially provides the LLM with relevant test failure information and then learns from both the failures and successes of previous patching attempts for the same bug, enhancing its ability for more effective APR (Xia and Zhang, 2023). GAMMA converts various fix templates into mask patterns and leverages a pre-trained language model to predict the correct code for the masked portions, treating APR as a fill-in-the-blank task (Zhang et al., 2023). Repilot generates a candidate patch by combining LLM suggestions with a Completion Engine, removing infeasible tokens and filling in gaps proactively (Wei et al., 2023). FitRepair integrates the plastic surgery hypothesis into LLM-based APR, combining the direct use of LLMs with two domain-specific fine-tuning strategies and one prompting strategy to enhance its repair capabilities. It can directly generate the correct code in context, effectively “filling in the blanks” of missing code lines or hunks (Xia et al., 2023a). Tare incorporated type checking into neural program repair model and can successfully repair 62 and 32 bugs from Defects4J v1.2 and Defects4J v2.0 (Zhu et al., 2023). GiantRepair creates patch skeletons from LLM-generated patches to narrow the patch space, then generates context-aware, high-quality patches by instantiating these skeletons for specific programs (Li et al., 2024). FixAgent unifies debugging through multi-agent collaboration and achieves strong performance with a three-layer hierarchical structure, where the final layer involves the use of a Web search engine (Lee et al., 2024). MORepair fine-tunes LLMs for program repair by adapting both to the syntactic nuances of code transformation and the underlying logic of code changes, enabling the generation of high-quality patches (Yang et al., 2024c). To leverage LLMs’ capabilities and augmented information, CREF is a semi-automatic repair framework for programming tutors, highlighting the potential for enhancing LLMs’ repair capabilities through tutor interactions and historical conversations (Yang et al., 2024b). RepairLLaMA is an innovative program repair method that finds optimal code representations for APR using fine-tuned models, and introduces a state-of-the-art parameter-efficient fine-tuning technique (PEFT) for program repair (Silva et al., 2023). Our approach differs from the aforementioned approaches in two key ways. First, we do not rely on fill-in-the-blank templates or skeletons for generating patches, which helps avoid patch overfitting problems in program repair (Smith et al., 2015; Le et al., 2018; Fei et al., 2024). Second, our fault localization step allows for the identification of multiple buggy methods, which are then fed into the LLM’s repair process along with various software artifacts. This provides flexibility to address not only single-method bugs but also bugs spanning across different methods. As seen in Table 7, the end-to-end process from localization to repair can generate plausible patches for 17.1% of non-single-method bugs, while most existing program repair approaches primarily target single-method bugs.
@@ 8. Conclusions and Future Work
This paper presents an LLM-based framework, DEVLoRe, for streamlining fault localization and program repair. By mimicking human developers in addressing bug problems and integrating three different software artifacts, DEVLoRe demonstrates strong performance in both fault localization and program repair, outperforming current state-of-the-art approaches in terms of bug fixing rate, time, and cost. In addition, unlike more rigid approaches that ask LLMs to fill in the blank within a single buggy method or use the fix template, DEVLoRe feed the LLM with only different software artifacts and there are no constraints on the DEVLoRe framework regarding how it repairs bugs, DEVLoRe has shown significant potential in handling bugs that span across multiple methods.
Our future work will focus on the following directions: first, expanding DEVLoRe to support additional programming languages, such as Python and C/C++, and evaluating its effectiveness on projects written in these languages; and second, testing the DEVLoRe framework with a broader range of software artifacts, such as commit history, code review comments, and user documentation, to assess how these additional data sources can further enhance the bug localization and program repair process.

### v3

@@ 6.3. Threats to Validity
Another potential threat is that the GPT-4o-mini model used in this work may have been trained on open-source projects from GitHub, which could overlap with the Defects4J dataset, leading to possible data leakage. To mitigate this, we also randomly selected 100 bugs from another dataset, GrowingBugs (Jiang et al., 2022), and found the plausible fixing rate to be 39%, which may help alleviate this concern. Additionally, the debugging information requires dynamic analysis, which is unlikely to have been used during the model’s training. Also as found in (Li et al., 2025; Campos et al., 2025), directly using the GPT-4o model can not improve the fix rate.
The final threat to external validity is that our systematic experiments used only one model and one approach, although using GPT-3.5 in DEVLoRe also outperform SOTA fault localization approaches. But the findings may not generalize to other models or approaches. We plan to explore this in our future work. Also, as discussed in Section 6.2, generalizing our framework to other programming languages, such as Python, does not currently achieve state-of-the-art performance. We plan to further improve and evaluate DEVLoRe on additional datasets across multiple programming languages in our future work.
@@ 7.1. Large Language Models for Fault Localization
Recently, there has been significant interest in using LLMs for fault localization. Toggle incorporated additional contextual information, such as the buggy line number or code review comments, and greatly enhances the accuracy of predicting both the starting and ending buggy tokens (Hossain et al., 2024). AGENTFL employs a multi-agent system based on ChatGPT and frames the fault localization task as a three-step process: comprehension, navigation, and confirmation. In each step, AGENTFL deploys specialized agents, each with unique expertise, and uses different tools to address specific tasks (Qin et al., 2025). CrashTracker conducts static analysis to map each crash to the corresponding exception instance and identify potential buggy candidates. It then utilizes LLMs to enhance the explainability of the localization results (Yan et al., 2024). Jiang et al. assessed the performance of recent commercial closed-source general-purpose LLMs, such as ChatGPT 3.5, ERNIE Bot 3.5, and IFlytek Spark 2.0, on line-level fault localization with the provided buggy method (Jiang et al., 2024). LLMAO fine-tunes LLMs with 350M, 6B, and 16B parameters on small, curated corpora like Defects4J, improving Top-1 fault localization by 2.3%-54.4% and Top-5 results by 14.4%-35.6%, compared to the state-of-the-art machine learning fault localization (Yang et al., 2024b). AutoFL prompts an LLM to use function calls for navigating a repository, enabling effective fault localization in large codebases while overcoming the LLM context length limit. It also generates an explanation of the bug and suggests a fault location (Kang et al., 2024). LLM4FL combines traditional spectrum-based fault localization with prompt chaining to divide large coverage data into manageable groups. By employing multiple LLM agents, it navigates the codebase more effectively to localize faults (Rafi et al., 2024). Like LLM4FL and AutoFL which combines transitional fault localization tools, DEVLoRe incorporates a static code analysis tool, MethodRecorder, to identify relevant buggy methods in a lightweight manner by invoking failing tests. However, unlike the above approaches, we provide the LLM with a combination of software artifacts, including issue content, error stack traces, and debugging information, which are commonly used by human developers for fault localization. This enables DEVLoRe to outperform state-of-the-art fault localization methods with greater efficiency and lower cost.
@@ 7.2. Large Language Models for Program Repair
Recent studies have extensively explored the use of LLMs for program repair. ChatRepair initially provides the LLM with relevant test failure information and then learns from both the failures and successes of previous patching attempts for the same bug, enhancing its ability for more effective APR (Xia and Zhang, 2023). GAMMA converts various fix templates into mask patterns and leverages a pre-trained language model to predict the correct code for the masked portions, treating APR as a fill-in-the-blank task (Zhang et al., 2023). Repilot generates a candidate patch by combining LLM suggestions with a Completion Engine, removing infeasible tokens and filling in gaps proactively (Wei et al., 2023). FitRepair integrates the plastic surgery hypothesis into LLM-based APR, combining the direct use of LLMs with two domain-specific fine-tuning strategies and one prompting strategy to enhance its repair capabilities. It can directly generate the correct code in context, effectively “filling in the blanks” of missing code lines or hunks (Xia et al., 2023a). Tare incorporated type checking into neural program repair model and can successfully repair 62 and 32 bugs from Defects4J v1.2 and Defects4J v2.0 (Zhu et al., 2023). GiantRepair creates patch skeletons from LLM-generated patches to narrow the patch space, then generates context-aware, high-quality patches by instantiating these skeletons for specific programs (Li et al., 2025). FixAgent unifies debugging through multi-agent collaboration and achieves strong performance with a three-layer hierarchical structure, where the final layer involves the use of a Web search engine (Lee et al., 2024). MORepair fine-tunes LLMs for program repair by adapting both to the syntactic nuances of code transformation and the underlying logic of code changes, enabling the generation of high-quality patches (Yang et al., 2024d). To leverage LLMs’ capabilities and augmented information, CREF is a semi-automatic repair framework for programming tutors, highlighting the potential for enhancing LLMs’ repair capabilities through tutor interactions and historical conversations (Yang et al., 2024c). RepairLLaMA is an innovative program repair method that finds optimal code representations for APR using fine-tuned models, and introduces a state-of-the-art parameter-efficient fine-tuning technique (PEFT) for program repair (Silva et al., 2023). Our approach differs from the aforementioned approaches in two key ways. First, we do not rely on fill-in-the-blank templates or skeletons for generating patches, which helps avoid patch overfitting problems in program repair (Smith et al., 2015; Le et al., 2018; Fei et al., 2025). Second, our fault localization step allows for the identification of multiple buggy methods, which are then fed into the LLM’s repair process along with various software artifacts. This provides flexibility to address not only single-method bugs but also bugs spanning across different methods. As seen in Table 7, the end-to-end process from localization to repair can generate plausible patches for 17.1% of non-single-method bugs, while most existing program repair approaches primarily target single-method bugs. A recent study (Parasaram et al., 2024) investigated how to identify an optimal set of facts to maximize the performance of LLM-based program repair. Their framework adopts chain-of-thought prompts, uses artifacts from both the fixed (correct) and buggy programs, and is evaluated on the BugsInPy dataset. Complementary to their work, we show that extracting all artifacts solely from the buggy version—which can be found in most development scenarios—and using direct prompts can still yield effective results on both the Defects4J dataset and SWE-bench Lite. Moreover, our paper demonstrates that different combinations of artifacts can enhance one another, helping LLMs on both fault localization and program repair.
@@ 8. Conclusions and Future Work
This paper presents an LLM-based framework, DEVLoRe, for streamlining fault localization and program repair. By mimicking human developers in addressing bug problems and integrating three different software artifacts, DEVLoRe demonstrates strong performance in both fault localization and program repair, outperforming current state-of-the-art approaches in terms of bug fixing rate, time, and cost on the Defects4J dataset. In addition, unlike more rigid approaches that ask LLMs to fill in the blank within a single buggy method or use the fix template, DEVLoRe feed the LLM with only different software artifacts and there are no constraints on the DEVLoRe framework regarding how it repairs bugs, DEVLoRe has shown significant potential in handling bugs that span across multiple methods.
Our future work will focus on the following directions: First, expanding DEVLoRe to support additional programming languages such as Python and C/C++, and evaluating and improving its effectiveness on projects written in these languages. Second, testing the DEVLoRe framework with a broader range of software artifacts—such as commit history, code review comments, and user documentation—to assess how these additional data sources can further enhance the bug localization and program repair process. Third, conducting a systematic evaluation involving multiple approaches and LLMs to better assess DEVLoRe’s generalization capabilities.